The sum of the first n terms of a sequence (Un) n € N* is given by Sn = n(n+1) / n+2. Find

a) The nth term of the sequence

b) The sum of the terms from the 6th to the 31st term inclusive and exclusive.



1
Expert's answer
2021-09-23T00:33:47-0400

a)


"S_n=S_{n-1}+a_n, n\geq2"

"a_n=S_n-S_{n-1}=\dfrac{n(n+1)}{n+2}-\dfrac{(n-1)(n-1+1)}{n-1+2}"

"=\dfrac{n(n^2+2n+1-n^2-2n+n+2)}{(n+1)(n+2)}"

"=\dfrac{n(n+3)}{(n+1)(n+2)}, n\geq2"

"a_1=S_1=\dfrac{1(1+1)}{1+2}=\dfrac{2}{3}"

b)


"S_5=\dfrac{5(5+1)}{5+2}=\dfrac{30}{7}"

"S_6=\dfrac{6(6+1)}{6+2}=\dfrac{21}{4}"

"S_{30}=\dfrac{30(30+1)}{30+2}=\dfrac{465}{16}"

"S_{31}=\dfrac{31(31+1)}{31+2}=\dfrac{992}{33}"

The sum of the terms from the 6th to the 31st term inclusive


"S_{31}-S_5=\dfrac{992}{33}-\dfrac{30}{7}=\dfrac{5954}{231}"

The sum of the terms from the 6th to the 31st term exclusive


"S_{30}-S_6=\dfrac{465}{16}-\dfrac{21}{4}=\dfrac{381}{16}"